Nine is subtracted from a number, and then the difference is multiplied by 5.
The result is 75. What is the number?

Found 2 solutions by mananth, robertb:
Answer by mananth(16946)

let number be x
9 is subtracted
x-9
multiplied by 5
5(x-9)
5(x-9)=75
5x-45=75
5x=45+75
5x=120
x= 60 the number
...
m.ananth@hotmail.ca

Answer by robertb(5830)